OpenStudy (anonymous):

How many solutions does the equation 5x + 10 = 5x − 8 have? Zero One Two Infinitely many

OpenStudy (anonymous):

@DIDITHEDAD

OpenStudy (anonymous):

Hold on...its been years since I have done this...give me a a minute as I go over review...

OpenStudy (mathmusician):

Do you know how to simplify the equation?

OpenStudy (anonymous):

kinda im new lol

OpenStudy (anonymous):

5x-5x=-8-10

OpenStudy (anonymous):

No solutions.... 5x - 5x = -8 - 10 because x=0 this means no solutions... AM I RIGHT GUYS? @MathMusician @joojoe

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

great!!!!!!

OpenStudy (anonymous):

Okay thanks...it is 0 because if x=0 then the answer will be zero.

OpenStudy (anonymous):

lol thank you... ;P all of you

OpenStudy (anonymous):

No problem....Have anymore questions?

OpenStudy (anonymous):

yep lol